The probability that a person picked at random is 85 years old is 0.008. If two people are picked at random, find the probability that:
a.) they are both 85
b.) neither is 85

Answer by ikleyn(52802)

.

(a)  P = 0.008*0.008.


(b)  P = (1-0.008)*(1-0.008) = 0.992*0.992.